# A NICER view of the â edge-on pulsar PSR J06143329

### v1

@@ Abstract
Four neutron star radius measurements have already been obtained by modeling the X-ray pulses of rotation-powered millisecond pulsars observed by the Neutron Star Interior Composition ExploreR (NICER). We report here the radius measurement of PSRÂ J06143329 employing the same method with NICER and XMM-Newton data using Bayesian Inference. For all different models tested, including one with unrestricted inclination prior, we retrieve very similar non-antipodal hot regions geometries and radii. For the preferred model, we infer an equatorial radius of km for a mass of (median values with equal-tailed credible interval), the latter being essentially constrained from radio timing priors obtained by MeerKAT. We find that, for all different models, the pulse emission originates from two hot regions, one at the pole and the other at the equator. The resulting radius constraint is consistent with previous X-ray and gravitational wave measurements of neutron stars in the same mass range. Equation of state inferences, including previous NICER and gravitational wave results, slightly soften the equation of state with PSRÂ J06143329 included and shift the allowed mass-radius region toward lower radii by m.
@@ 1 Introduction
Neutron stars (NSs) are known to harbor ultra-dense matter in their cores (up to ), whose microscopic state can be constrained by the macroscopic properties of NSs. The Neutron Star Interior Composition ExploreR (NICER, Gendreau et al. 2016) is a fast X-ray timing mission, whose primary goal is the measurement of the radii of fast rotating NSs, known as millisecond pulsars (MSPs), to yield constraints on their inside matter.
@@
MSPs are produced through a recycling process, where a NS accretes matter from a companion star, inducing its spin-up to millisecond rotation periods via transfer of angular momentum (Alpar et al., 1982). Such accelerated MSPs have a magnetic field strong enough (G) to cause the acceleration of charged particles in the magnetosphere, which results in complex and pulsed multi-wavelength emission as the NS rotates. In particular, the magnetic field induces an electron-positron particle shower near the magnetic poles of the NS surface . This phenomenon deposits energy at the magnetic polar caps of the NS, which locally heats the surface to K, consequently emitting thermal radiation in the soft X-ray (Ruderman & Sutherland, 1975; Arons, 1981; Harding & Muslimov, 2001, 2002) observed as pulsed emission as the NS rotates.
@@
Measurements of MSPs radii are carried out by estimating their compactness through the imprint of their strong gravitational field on the pulse profile from the aforementioned surface emission. Contrary to other kinds of NS emission (e.g., radio, non-thermal hard X-rays or -rays), such surface emission is strongly affected by spacetime curvature due to the vicinity of the emission region to the center of the gravitational potential (km). Other relativistic effects, such as Doppler boosting and angle aberration, also carry information on the radius. In practice, pulse profile modeling (PPM) is used to model these effects and to infer the source parameters, including its radius, and their uncertainties using statistical methods (Pavlov & Zavlin, 1997; Zavlin & Pavlov, 1998; Bogdanov, 2013; Watts, 2019; Bogdanov et al., 2021). This method has now been extensively tested and used on four different MSPs observed by NICER, with different datasets (Riley et al., 2019, 2021; Miller et al., 2019, 2021; Salmi et al., 2022, 2023; Choudhury et al., 2024a; Dittmann et al., 2024; Salmi et al., 2024a, b; Vinciguerra et al., 2024; Hoogkamer et al., 2025). Such methods enable precise modeling of the surface emission at the cost of high computation requirements.
@@
Mass and radius measurements of NSs lead to constraints on the equation of state (EOS) of dense nuclear matter by informing about its stiffness (Watts et al., 2016; Chatziioannou et al., 2024). Around nuclear densities, the EOS is informed by chiral effective field theory (EFT, see, e.g., Hebeler et al. 2013; Tews et al. 2013; Drischler et al. 2019; Keller et al. 2023; Tews et al. 2024). The electromagnetic wave measurements obtained with PPM, along with gravitational wave (GW) measurements (GW170817, Abbott et al. 2018; GW190425, Abbott et al. 2020b), currently favor EOSs on the stiffer side of a EFT informed prior distribution (e.g., Raaijmakers et al. 2021; Rutherford et al. 2024; Koehn et al. 2025). While nucleonic EOS are the simplest explanation, more exotic EOS families also exist (see Lattimer & Prakash 2016; Oertel et al. 2017; Burgio et al. 2021 and references therein for examples). However, none of these EOS models are significantly favored by observations, which highlights the need for more numerous and tighter observational constraints.
@@
The MSP analyzed in this study is PSR J06143329, an intermediate mass MSP rotating at Hz discovered from its radio pulsations (Ransom et al., 2011). It has also been found to exhibit soft X-ray pulsations, likely of thermal origin, with no sign of non-thermal pulsations as all the pulsed signal was below keV (Guillot et al., 2019). A white dwarf (WD) companion with a hydrogen atmosphere has been discovered (Bassa et al., 2016), allowing for distance estimates using the photometric temperature of the companion. This distance was later confirmed with radio parallax (Shamohammadi et al., 2024). Tight constraints on the mass and orbital inclination also exist from strong Shapiro delay caused by the edge-on view of the PSR J06143329 system (Reardon et al., 2023; Miles et al., 2025). Such prior knowledge greatly decreases possible degeneracies with PPM and provide important insights to yield a tight and robust radius measurement, which proves more complicated without prior mass and inclination information (Vinciguerra et al., 2024; Salmi et al., 2024b). PSR J06143329 has been observed by both NICER and the X-ray Multi-Mirror Mission (XMM-Newton, XMM hereafter), with a deep NICER accumulated exposure of Ms. It is hence a good candidate for a new mass-radius measurement using PPM.
@@ 2 X-ray Event data
This section summarizes the reduction of the X-ray data used for inference. All the data products used for this analysis can be found in the Zenodo repository111https://doi.org/10.5281/zenodo.15603406.
@@ 2.1 NICER
This analysis of PSRÂ J06143329 makes use of 1.6âMs of raw NICER data acquired in the 2018 March 3rd â 2023 May 15 period (OBSIDs to ), comprised of all the observations of this source antecedent to the NICER light leak222https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/light-leak-overview/. The few available observations dating after the light leak occurrence (19âks) are characterized by a high noise level caused by optical loading and were not included due to the difficulty of properly screening them alongside their short additional exposure.
The screening procedure, to remove time intervals with high background, generally follows the one adopted in Bogdanov et al. (2019) and by the other NICER PPM analyses listed above. Calibration was handled by the Heasoft software (Nasa High Energy Astrophysics Science Archive Research Center (2014), HEASARC), version 6.33.2, which includes NICERDASv12 along with the calibration file xti20240206. Initial screening was produced with the nicerl2 task. Unlike previous analyses, no âhotâ focal plane modules (FPMs) were consistently screened out since, as part of the niautoscreen task, they are now individually removed when exhibiting count rates above the median of active FPMs. The final ancillary response file (ARF) is automatically scaled down accordingly, based on the average number of FPMs activated during the whole exposure. This allows to extract more events while limiting background.
Good time intervals (GTIs) were further selected with the psrpipe script from NICERsoft333https://github.com/paulray/NICERsoft, by removing time intervals which included any of the following:
@@
High optical photon flux causing considerable FPMs reset: Median Undershoot rate cts/s.
@@
Extracted events were phase-folded with the PINT task photonphase (Luo et al., 2021), using a joint Nançay Radio Telescope (NRT) and Fermi timing solution spanning the entire NICER observation window, to produce a phase-resolved spectrum (Figure 1). The resulting bolometric pulse profile has a strong main peak and a weaker secondary peak, which is consistent with previous findings (Guillot et al., 2019). Finally, the associated ARF and response matrix file (RMF) were extracted using nicerl3-spect task.
@@
The analyses were performed on pulse invariant (PI) energy channels , corresponding to an energy band of keV. While the lower energy bound is motivated by the NICER effective area sudden drop as well as residual optical photon flux background below keV, the upper bound is motivated by the low source flux and absence of pulsed emission over keV, making the total flux background dominated above this energy. A Kolmogorov-Smirnov test comparing the phase distribution of photons with energy over keV to a uniform distribution did not find any significant deviation (). All inferences results pointed toward a contribution of the source to the total NICER flux lower than at keV (see Section 4), validating the choice of this upper limit.
@@
The NICER phase-averaged spectrum shows two prominent features near keV and keV respectively (see Section 4.4). Both are phase-independent peaks, the latter appearing mostly in observations after mid-2022. We identify it as a fluorescence line of oxygen VII at eV in the Earth atmosphere, which is excited by the stronger solar activity at later dates. This line also already appeared prominently in NICER spectra of some analyses (Salmi et al., 2024a, b) and subtly for other analyses that did not include any observations after the solar activity started rising (mid-2022), and for which the source flux was higher (Choudhury et al., 2024a; Vinciguerra et al., 2024).
@@
MSP surface emission models cannot explain such time-independent narrow spectral features. Hence, such an excess is accounted for as a phase-independent background, which is fitted alongside the modeled pulse profile (more details in Section 3.5). We ensured that these spectral features were well reproduced by the fitted phase-independent background by inspecting the inferred background for all runs. Such peaks being also inferred as background in previous analyses further supports that PPM accounts for them in the phase-independent background, even in cases where the O VII line was less prominent (Vinciguerra et al., 2024).
@@
Although NICER background models such as Space Weather (SW) or 3C50 (Remillard et al., 2022) exist, none has been used in this study and the NICER background is inferred directly from the data. This approach is facilitated by the available XMM observations which informs on the phase-averaged source spectrum and flux, which in turn informs the NICER expected source spectrum, and hence its background spectrum. Salmi et al. (2022) has shown that including background constraints from XMM for PSR J07406620 gives comparable results to using SW or 3C50 without relying on any of the assumptions that either of them makes. Moreover, 3C50 relies on observations from before mid-2020, when the O VII line was not as prominent, which can lead to underestimation of the contribution of this background component to the flux. As one XMM observation was available, we decided to use it instead of one of the background models.
@@ 2.2 XMM-Newton
The source spectra were extracted from a 35âłÂ circular region, while the backgrounds were extracted from a large 35-550âłÂ annulus region centered on the source, and limited to the same charge-coupled device. Other X-ray sources were excluded and low vignetting regions avoided. The exact regions differ between EPIC instruments. The backscal parameter and respective exposure times written in the processed observation files were used to scale the backgrounds spectra to the sources spectra. The ARF and RMF were extracted using the arfgen and rmfgen tasks.
@@ 3 Modeling
The modeling procedure as a whole is similar to that of Riley et al. (2021); Salmi et al. (2022); Choudhury et al. (2024a); Salmi et al. (2024a); Vinciguerra et al. (2024). This section summarizes the PPM, the parameter inference method and adopted settings. All relevant files for modeling and inference are provided in the Zenodo repository.
@@ 3.1 X-PSI settings
We model the energy dependent pulse profile using the X-ray Pulse Simulation and Inference package (X-PSI, Riley et al. 2023) v2.2.7. For post-processing of the inference runs only, we used X-PSI v3.1.0.
The resolution settings of X-PSI (see Appendix A for explanations on the different settings) have below 1% effects on the computed waveform for simple geometries and can reach a few percent under extreme cases with thin and elongated spot structures (Choudhury et al., 2024b). Due to the limited count statistics of our dataset, we deemed such complex geometries to be difficult to distinguish from simpler ones. Hence, these were not investigated during our exploration runs which were performed using the low resolution (LR) X-PSI settings from Choudhury et al. (2024b). This allowed a lower computational cost and the use of a better sampling resolution instead. However, our headline result uses higher X-PSI resolution to compute accurate waveforms for those extreme cases and mitigate the effects of a limited X-PSI resolution on the sampling procedure (see Appendix A). These effects are further investigated in Section 5.
@@ 3.2 Sampling
The posterior distributions of model parameters were obtained with nested sampling (Skilling, 2004). Among other Bayesian inference algorithms, nested sampling has the advantage of providing thorough sampling over a large prior parameter space and easy model comparison with limited number of hyper-parameters. In practice, sampling was carried out using PyMultiNest (Buchner et al., 2014), a python wrapper of the nested sampling implementation MultiNest (Feroz et al., 2009). Although UltraNest (Buchner, 2021) is in theory more robust, it is also much slower and performed equally well as MultiNest for PPM on PSR J07406620 (Hoogkamer et al., 2025). We hence opted to use MultiNest for its lower computational cost.
All exploratory runs referenced here used the same set of hyper-parameters, chosen to ensure good sampling of the parameter space, while limiting computation time and allowing for easy and fair comparison between runs. MultiNest has three hyper-parameters, which were chosen to be live points (LP), a sampling efficiency (SE) of and an evidence tolerance (ET) of for the exploratory runs (see Feroz et al. 2009 for description of these hyper-parameters). We ran all inferences with multi-modal exploration enabled. This allocates distinct sampling regions to the different modes identified, allow for individual post processing of the latter, and may speed-up the computations (Feroz et al., 2009). Such hyper-parameters grant a more complete sampling (for exploratory runs) than in previous studies. Our headline result was run with LP, SE and ET, which allows for a higher sampling accuracy. The exploration of the parameter space completeness is further discussed in Section 5.

### v2

# A NICER view of the â edge-on pulsar PSRÂ J06143329

@@ Abstract
Four neutron star radius measurements have already been obtained by modeling the X-ray pulses of rotation-powered millisecond pulsars observed by the Neutron Star Interior Composition ExploreR (NICER). We report here the radius measurement of PSRÂ J06143329 employing the same method with NICER and XMM-Newton data using Bayesian inference. For all different models tested, including one with unrestricted inclination prior, we retrieve very similar non-antipodal hot region geometries and radii. For the preferred model, we infer an equatorial radius of km for a mass of (median values with equal-tailed credible interval), the latter being essentially constrained from radio timing priors obtained by MeerKAT. A more complex model, fitting the data equally well, resulted in a consistent inferred radius. We find that, for all different models, the pulse emission originates from two hot regions, one at the pole and the other at the equator. The resulting radius constraint is consistent with previous X-ray and gravitational wave measurements of neutron stars in the same mass range. Equation of state inferences, including previous NICER and gravitational wave results, slightly soften the equation of state with PSRÂ J06143329 included and shift the allowed mass-radius region toward lower radii by m, which is compatible with previous analyses to within less than one standard deviation.
@@ 1 Introduction
Neutron stars (NSs) are known to harbor ultra-dense matter in their cores (up to ), whose microscopic state can be constrained by the macroscopic properties of NSs. The Neutron Star Interior Composition ExploreR (NICER, gendreau_neutron_2016) is a fast X-ray timing mission, whose primary goal is the measurement of the radii of fast rotating NSs, known as millisecond pulsars (MSPs), to yield constraints on their inside matter.
@@
MSPs are produced through a recycling process, where an NS accretes matter from a companion star, inducing its spin-up to millisecond rotation periods via transfer of angular momentum (alpar_new_1982). Such accelerated MSPs have a magnetic field strong enough (G) to cause the acceleration of charged particles in the magnetosphere, which results in complex and pulsed multi-wavelength emission as the NS rotates. In particular, the magnetic field induces an electron-positron particle shower near the magnetic poles of the NS surface. This phenomenon deposits energy at the magnetic polar caps of the NS, which locally heats the surface to K, consequently emitting thermal radiation in the soft X-ray (ruderman_theory_1975; arons_pair_1981; harding_pulsar_2001; harding_pulsar_2002) observed as pulsed emission as the NS rotates.
@@
Measurements of MSPs radii are carried out by estimating their compactness through the imprint of their strong gravitational field on the pulse profile from the aforementioned surface emission. Contrary to other kinds of NS emission (e.g., radio, non-thermal hard X-rays or -rays), such surface emission is strongly affected by spacetime curvature due to the vicinity of the emission region to the center of the gravitational potential (km). Other relativistic effects, such as Doppler boosting and angle aberration, also carry information on the radius. In practice, pulse profile modeling (PPM) is used to model these effects and to infer the source parameters, including its radius, and their uncertainties using statistical methods (pavlov_mass--radius_1997; zavlin_soft_1997; bogdanov_nearest_2013; watts_constraining_2019; bogdanov_constraining_2021). This method has now been extensively tested and used on four different MSPs observed by NICER, with different datasets (riley_nicer_2019; riley_nicer_2021; miller_psr_2019; miller_radius_2021; salmi_radius_2022; salmi_atmospheric_2023; choudhury_nicer_2024; dittmann_more_2024; salmi_radius_2024; salmi_J1231_2024; vinciguerra_updated_2024; hoogkamer_cross-comparison_2025). Such methods enable precise modeling of the surface emission at the cost of high computation requirements.
@@
Mass and radius measurements of NSs lead to constraints on the equation of state (EOS) of dense nuclear matter by informing about its stiffness (Watts:2016uzu; Chatziioannou:2024tjq). Around nuclear densities, the EOS is informed by chiral effective field theory (EFT, see, e.g., Hebeler2013; Tews:2012fj; drischler_chiral_2019; Keller2023; Tews:2024owl). The electromagnetic wave measurements obtained with PPM, along with gravitational wave (GW) measurements (GW170817, abbott_gw170817_2018; GW190425, abbott_gw190425_2020), currently favor EOSs on the stiffer side of a EFT informed prior distribution (e.g., Raaijmakers21; rutherford_constraining_2024; Koehn:2024set). While nucleonic EOS are the simplest explanation, more exotic EOS families also exist (see lattimer_equation_2016; oertel_equations_2017; burgio_neutron_2021 and references therein for examples). However, none of these EOS models are significantly favored by observations, which highlights the need for more numerous and tighter observational constraints.
@@
The MSP analyzed in this study is PSRÂ J06143329, an intermediate mass MSP rotating at Hz discovered from its radio pulsations (ransom_three_2011). It has also been found to exhibit soft X-ray pulsations, likely of thermal origin, with no sign of non-thermal pulsations as all the pulsed signal was below keV (guillot_nicer_2019). A white dwarf (WD) companion with a hydrogen atmosphere has been discovered (bassa_cool_2016), allowing for distance estimates using the photometric temperature of the companion. This distance was later confirmed with radio parallax (shamohammadi_meerkat_2024). Tight constraints on the mass and orbital inclination also exist from strong Shapiro delay caused by the edge-on view of the PSRÂ J06143329 system (reardon_gravitational-wave_2023; miles_meerkat_2025). Such prior knowledge greatly decreases possible degeneracies with PPM and provide important insights to yield a tight and robust radius measurement, which proves more complicated without prior mass and inclination information (vinciguerra_updated_2024; salmi_J1231_2024). PSRÂ J06143329 has been observed by both NICER and the X-ray Multi-Mirror Mission (XMM-Newton, XMM hereafter), with a deep NICER accumulated exposure of Ms. It is hence a good candidate for a new mass-radius measurement using PPM.
@@ 2 X-Ray Event data
This section summarizes the reduction of the X-ray data used for inference. All the data products used for this analysis can be found in the Zenodo repository doi: 10.5281/zenodo.15603405 (zenodo).
@@ 2.1 NICER
This analysis of PSRÂ J06143329 makes use of 1.6âMs of raw NICER data acquired in the 2018 March 3rd â 2023 May 15 period (OBSIDs to ), comprised of all the observations of this source antecedent to the NICER light leak111https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/light-leak-overview/. The few available observations dating after the light leak occurrence (19âks) are characterized by a high noise level caused by optical loading and were not included due to the difficulty of properly screening them alongside their short additional exposure.
The screening procedure, to remove time intervals with high background, generally follows the one adopted in Bogdanov2019a and by the other NICER PPM analyses listed above. Calibration was handled by the Heasoft software (2014ascl.soft08004N), version 6.33.2, which includes NICERDASv12 along with the calibration file xti20240206. Initial screening was produced with the nicerl2 task. Unlike previous analyses, no âhotâ focal plane modules (FPMs) were consistently screened out since, as part of the niautoscreen task, they are now individually removed when exhibiting count rates above the median of active FPMs. The final ancillary response file (ARF) is automatically scaled down accordingly, based on the average number of FPMs activated during the whole exposure. This allows to extract more events while limiting background.
Good time intervals (GTIs) were further selected with the psrpipe script from NICERsoft222https://github.com/paulray/NICERsoft, by removing time intervals which included any of the following:
@@
High optical photon flux causing considerable FPM reset: median undershoot rate cts/s.
@@
Extracted events were phase-folded with the PINT task photonphase (luo_pint_2021), using a joint Nançay Radio Telescope (NRT) and Fermi timing solution spanning the entire NICER observation window, to produce a phase-resolved spectrum (Figure 1). The resulting bolometric pulse profile has a strong main peak and a weaker secondary peak, which is consistent with previous findings (guillot_nicer_2019). Finally, the associated ARF and response matrix file (RMF) were extracted using nicerl3-spect task.
@@
The analyses were performed on pulse invariant (PI) energy channels , corresponding to an energy band of keV. While the lower energy bound is motivated by the NICER effective area sudden drop as well as residual optical photon flux background below keV, the upper bound is motivated by the low source flux and absence of pulsed emission over keV, making the total flux background dominated above this energy. A Kolmogorov-Smirnov test comparing the phase distribution of photons with energy over keV to a uniform distribution did not find any significant deviation (). All inferences pointed toward a contribution of the source to the total NICER flux lower than at keV (see Section 4), validating the choice of this upper limit.
@@
The NICER phase-averaged spectrum shows two prominent features near keV and keV respectively (see Section 4.4). Both are phase-independent peaks, the latter appearing mostly in observations after mid-2022. We identify it as a fluorescence line of oxygen VII at eV in the Earth atmosphere, which is excited by the stronger solar activity at later dates. This line also already appeared prominently in NICER spectra of some analyses (salmi_radius_2024; salmi_J1231_2024) and subtly for other analyses that did not include any observations after the solar activity started rising (mid-2022), and for which the source flux was higher (choudhury_nicer_2024; vinciguerra_updated_2024).
@@
MSP surface emission models cannot explain such time-independent narrow spectral features. Hence, such an excess is accounted for as a phase-independent background, which is fitted alongside the modeled pulse profile (more details in Section 3.5). We ensured that these spectral features were well reproduced by the fitted phase-independent background by inspecting the inferred background for all runs. Such peaks being also inferred as background in previous analyses further supports that PPM accounts for them in the phase-independent background, even in cases where the O VII line was less prominent (vinciguerra_updated_2024).
@@
Although NICER background models such as Space Weather (SW) or 3C50 (remillard_empirical_2022) exist, none has been used in this study and the NICER background is inferred directly from the data. This approach is facilitated by the available XMM observations, which inform on the phase-averaged source spectrum and flux, which in turn inform the NICER expected source spectrum, and hence its background spectrum. salmi_radius_2022 has shown that including background constraints from XMM for PSRÂ J07406620 gives comparable results to using SW or 3C50 without relying on any of the assumptions that either of them makes. Moreover, 3C50 relies on observations from before mid-2020, when the OÂ VII line was not as prominent, which can lead to underestimation of the contribution of this background component to the flux. As one XMM observation was available, we decided to use it instead of one of the background models.
@@ 2.2 XMM-Newton
The source spectra were extracted from a 35âł circular region, while the backgrounds were extracted from a large 35-550âł annulus region centered on the source, and limited to the same charge-coupled device. Other X-ray sources were excluded and low vignetting regions avoided. The exact regions differ between EPIC instruments. The backscal parameter and respective exposure times written in the processed observation files were used to scale the background spectra to the sources spectra. The ARF and RMF were extracted using the arfgen and rmfgen tasks.
@@ 3 Modeling
The modeling procedure as a whole is similar to that of riley_nicer_2021; salmi_radius_2022; choudhury_nicer_2024; salmi_radius_2024; vinciguerra_updated_2024. This section summarizes the PPM, the parameter inference method and adopted settings. All relevant files for modeling and inference are provided in the Zenodo repository (zenodo).
@@ 3.1 X-PSI settings
We model the energy dependent pulse profile using the X-ray Pulse Simulation and Inference package (X-PSI, xpsi) v2.2.7. For post-processing of the inference runs only, we used X-PSI v3.1.0.
The resolution settings of X-PSI (see Appendix A for explanations on the different settings) have below 1% effects on the computed waveform for simple geometries and can reach a few percent under extreme cases with thin and elongated spot structures (choudhury_exploring_2024). Due to the limited count statistics of our dataset, we deemed such complex geometries to be difficult to distinguish from simpler ones. Hence, these were not investigated during our exploration runs which were performed using the low resolution (LR) X-PSI settings from choudhury_exploring_2024. This allowed a lower computational cost and the use of a better sampling resolution instead. However, our headline result uses higher X-PSI resolution to compute accurate waveforms for those extreme cases and mitigate the effects of a limited X-PSI resolution on the sampling procedure (see Appendix A). These effects are further investigated in Section 5.
@@ 3.2 Sampling
The posterior distributions of model parameters were obtained with nested sampling (skilling2004nested). Among other Bayesian inference algorithms, nested sampling has the advantage of providing thorough sampling over a large prior parameter space and easy model comparison with limited number of hyper-parameters. In practice, sampling was carried out using PyMultiNest (PyMultiNest), a Python wrapper of the nested sampling implementation MultiNest (MultiNest_2009). Although UltraNest (buchner_ultranest_2021) is in theory more robust, it is also much slower and performed equally well as MultiNest for PPM on PSRÂ J07406620 (hoogkamer_cross-comparison_2025). We hence opted to use MultiNest for its lower computational cost.
All exploratory runs referenced here used the same set of hyper-parameters, chosen to ensure good sampling of the parameter space, while limiting computation time and allowing for easy and fair comparison between runs. MultiNest has three hyper-parameters, which were chosen to be live points (LP), a sampling efficiency (SE) of and an evidence tolerance (ET) of for the exploratory runs (see MultiNest_2009 for a description of these hyper-parameters). We ran all inferences with multi-modal exploration enabled. This allocates distinct sampling regions to the different modes identified, allows for individual post processing of the latter, and may speed-up the computations (MultiNest_2009). Such hyper-parameters grant a more complete sampling (for exploratory runs) than in previous studies. Our headline result was run with LP, SE and ET, which allows for a higher sampling accuracy. The exploration of the parameter space completeness is further discussed in Section 5.